OpenStudy (stamp):

Let Ø be an angle in the quadrant III such that sinø= -3/5. Find the exact values of secø and cotø.

OpenStudy (stamp):

secø = -5/3 and cotø = 4/3?

OpenStudy (stamp):

jk let me try this again

OpenStudy (vishweshshrimali5):

Signs are correct :) cot (phi) is correct Check sec(phi) again

OpenStudy (stamp):

secø = -4/5 cotø = 3/4

OpenStudy (vishweshshrimali5):

Why is cot (phi) = 3/4 ?

OpenStudy (stamp):

omg secø = -4/5 cotø = 4/3

OpenStudy (stamp):

tan = 3/4 i forgot to flip

OpenStudy (vishweshshrimali5):

Yup :) Now they are correct

OpenStudy (stamp):

9999%??

OpenStudy (vishweshshrimali5):

+1 %

OpenStudy (stamp):

grazie

OpenStudy (vishweshshrimali5):

:D Great work :)